# Empirical Best Prediction of Poverty Indicators via Nested Error Regression with High Dimensional Parameters

### v1

@@ 6.2 Experiment 2: 17 out-of-sample provinces
Table 3: Performance of estimators/predictors of small area FGT measures when some areas are not sampled; Average is over 17 out-of-sample areas; numbers in parenthesis are the values of the efficiency over MR in terms of RRMSPE.
Predictor RRMSPE % (EFF) Relative Bias %
Headcount Ratio Poverty Gap Headcount Ratio Poverty Gap
MR 35.36 (1.00) 41.87 (1.00) 31.50 29.78
CLS 27.49 (0.78) 16.69 (0.40) 18.83 -0.93
ELL 26.04 (0.74) 20.55 (0.49) 19.17 10.46
Figure 4(a) illustrates the province-specific performance for the 17 out-of-sample provinces, demonstrating that the CLS model consistently outperforms both MR and ELL for the poverty gap measure in terms of RRMSPE and relative bias. Notably, CLS achieves significantly smaller RRMSPE than MR for the majority of out-of-sample areas. This superior performance can be explained by the flexibility of the CLS model, which allows regression coefficients and sampling variances to vary across areas, enabling it to better capture region-specific heterogeneity.
For the headcount ratio, Figure 4(b) shows that CLS also performs better than MR, although the margin of improvement is smaller compared to the poverty gap measure. These results highlight the capability of the CLS model to adapt to area-level differences, offering enhanced precision and reduced bias in poverty estimates.
(a) The RRMSPE and Relative Bias of the EBP estimators of poverty gap for out of sample areas.
(b) The RRMSPE and Relative Bias of the EBP estimators of headcount ratio for out of sample areas.
Figure 4: The RRMSPE and Relative Bias of the EBP estimators of FGT measures for out of sample areas.
@@ 7 Application: poverty mapping for Albania
Poverty mapping is a crucial tool for understanding the spatial distribution of poverty and guiding the implementation of poverty alleviation programs. In Albania, where the economy is predominantly rural and income is difficult to measure accurately, poverty is assessed using a consumption-based measure (Betti, Neri and Bellini, 2003). In this section, we use 2002 Living Standards Measurement Survey (LSMS) data to estimate poverty indicators for 374 municipalities in Albania. To supplement this information, we incorporate auxiliary data from the 2001 Census, which covers 726,895 households across Albania. The descriptive statistics of these datasets have been presented in Section 3, providing an overview of their key characteristics and ensuring their suitability for small area estimation.”
Design-based simulation results in Section 6 indicate that the CLS method is a promising approach due to: (i) significantly lower bias and RRMSPE compared to the MR method and (ii) greater computational efficiency relative to the existing algorithm proposed in Lahiri and Salvati (2023), improving feasibility for large-scale datasets.
@@
To estimate poverty indicators for Albanian municipalities, we consider household consumption expenditure as the primary welfare variable, denoted as $E_{ij}$. Following Section 6, we apply a shifted logarithm transformation to obtain $Y_{ij}$ as the final response variable in the NERHDP model. The covariates, including household size, dwelling facilities, and ownership variables, are based on prior poverty assessment studies in Albania (Betti, Neri and Bellini, 2003; Tzavidis et al., 2008) and have been presented in Section 3. The poverty line is set at 4,891 Leks per month for all districts, based on Betti, Neri and Bellini (2003).
For each small area, we estimate the sample variance and fit a regression model where the response variable is the shifted logarithm transformed household consumption expenditure, with covariates previously specified from the LSMS data. Initial data exploration suggests substantial heterogeneity in both within-area variation and regression coefficients across municipalities. Figure 5 presents box plots of the distributions of estimated regression coefficients (top panel) and residual distributions by municipality (bottom panel). These results indicate that assuming identical regression coefficients and/or variance components in the MR method is unrealistic for this dataset, emphasizing the need for a more flexible modeling approach that accounts for area-specific variations.
@@
(a) Distributions of estimated regression coefficients.
(b) Distributions of the residuals by municipality.
Figure 5: Distributions of estimated regression coefficients (a) and residuals (b) obtained by fitting a regression model separately for each municipality.
@@
We apply the CLS method to estimate empirical best predictors (EBPs) of headcount ratios and poverty gaps by municipality, denoted as $\hat{F}_{0i}^{CLS}$ and $\hat{F}_{1i}^{CLS}$, respectively. The MSPEs of these FGT measures are estimated using the parametric bootstrap estimator $\text{mspe}_{*}(\hat{F}_{\alpha i}^{CLS})$ with $B=100$ replicates. We then compute the associated coefficients of variation (CVs) using:
@@
Following Molina and Rao (2010), we also compute direct estimators of the FGT measures $\hat{F}_{\alpha i}^{w}$ and their variance estimates $\widehat{\text{Var}}(\hat{F}_{\alpha i}^{w})$ for sampled municipalities using weighted sample means (see Section 3).
To assess the quality of our CLS estimator, we use a set of diagnostic tools based on the requirement that model-based small area estimates should align with the corresponding unbiased direct estimates, albeit more precise (Marino et al., 2019). Correlation coefficients between CLS and direct estimates for the headcount ratio and poverty gap are 0.796 and 0.802, respectively, suggesting strong agreement between the two approaches. We further evaluate coherence using the goodness-of-fit diagnostic (Brown et al., 2001):
@@
With $W=119.59$ for headcount ratio and $W=175.75$ for poverty gap, both are below the 95th percentile of a $\chi^{2}$ distribution with 212 degrees of freedom ($\chi^{2}_{212,0.95}=177.39$), indicating that CLS estimates are coherent with direct estimates.
To assess precision gains, Table 4 presents summary statistics for CLS and direct estimates across municipalities. The results indicate that CLS estimators consistently exhibit lower CVs than direct estimators, even in municipalities with varying sample sizes. Moreover, for some municipalities where the estimated headcount ratio and poverty gap are zero, the CVs of direct estimates are unavailable, highlighting a key limitation of the direct estimation method.
Table 4: Population size, sample size, direct and CLS estimates of headcount ratios (%) and poverty gaps (%), and the associated CVs (%) of direct and CLS estimators for the Albania municipalities with sample size closest to minimum, first quartile, median, third quartile, and maximum.
Headcount Ratio Poverty Gap Municipality $N_{i}$ $n_{i}$ $\hat{F}_{0i}^{w}$ $\hat{F}_{0i}^{CLS}$ CV($\hat{F}_{0i}^{w}$) CV($\hat{F}_{0i}^{CLS}$) $\hat{F}_{1i}^{w}$ $\hat{F}_{1i}^{CLS}$ CV($\hat{F}_{1i}^{w}$) CV($\hat{F}_{1i}^{CLS}$) Hajmel $1$ 6 33.33 23.82 56.96 27.01 1.52 5.92 70.56 31.31 Orenje 1419 16 12.50 20.69 66.19 15.06 2.00 4.97 66.03 17.44 B.Curri 1488 40 20.00 24.54 32.82 8.57 5.65 5.51 43.01 9.56 Gramsh 2503 64 23.55 17.97 24.48 0.69 4.75 3.64 38.93 21.90 Tirane 89764 600 12.43 11.72 10.91 4.46 2.49 2.31 14.71 4.72
We further compare the empirical cumulative density functions (ECDFs) of the CV of the CLS estimator and the direct estimator. Figure 6 presents CV (%) values from all sampled municipalities and shows that the ECDFs for the CLS estimators, for both the headcount ratio and the poverty gap, consistently dominate those of the direct estimators, further highlighting the improved precision of CLS estimators. Additionally, we find that about $78\%$ and $89\%$ of Albanian municipalities have CV values associated with the direct estimates of the headcount ratio and poverty gap, respectively, exceeding the conventional 33% reliability threshold in small area estimation (Marino et al., 2019). In contrast, these percentages reduce to 28% and 34% when applying the proposed CLS approach.
@@
We generate poverty maps for Albania using the direct estimation method, as shown in Figure 7. The gray areas represent regions where poverty estimates are unavailable due to insufficient sample sizes, underscoring a key limitation of the direct estimation approach in small-area analysis. In contrast, Figure 8 presents poverty maps generated using our proposed CLS method, which provides more comprehensive and reliable estimates across all municipalities. These maps reveal that several municipalities in the Bulqize district—such as Ostren, Gjorice, Shupenze, and Bulqize — exhibit the highest headcount ratio and poverty gap values, significantly exceeding those observed in other municipalities. Similarly, municipalities like Gjoaaj in the Peqin district, Katund I in the Durres district, and Arren in the Kukes district also show relatively high poverty levels, highlighting the need for targeted policy interventions to improve living conditions in these areas. In contrast, municipalities in the southern regions display substantially lower poverty levels, reflecting better economic conditions and higher living standards.
Figure 7: Municipality-level direct estimates of headcount ratios and poverty gaps in Albania.
Figure 8: Municipality-level CLS estimates of headcount ratios and poverty gaps in Albania.
These results are consistent with the findings of Betti, Neri and Bellini (2003), who used the ELL method, and those of Tzavidis et al. (2008), further validating the robustness of our approach in capturing the spatial distribution of poverty in Albania.
The CLS method offers a significant advantage over direct estimation by producing estimates for all municipalities, thereby addressing data gaps observed in the direct method. This improvement ensures a more complete understanding of poverty distribution, which is crucial for effective policy planning.
The identification of high-poverty areas, particularly in the northern and central regions, highlights the need for geographically targeted poverty alleviation programs. In contrast, the relatively lower poverty levels in the south suggest regional disparities that should be accounted for in national development strategies.
The consistency of CLS estimates across municipalities, even in those with limited survey data, reinforces the reliability of this approach in small area poverty estimation. This robustness makes it a valuable tool for informing social protection policies and resource allocation.
@@ 8 Discussion
The proposed model bridges the gap between existing approaches that either assume identity in model parameters across areas, potentially leading to misspecification, or rely on random effect specifications that demand strong distributional assumptions. By leveraging area-specific estimating equations and appropriately constructed residuals, we offer a framework for deriving empirical best prediction of complex poverty measures that captures variability across areas without the instability associated with fixed effects in small samples. On the computational side, we introduce a novel, computational efficient, and data-driven procedure for estimating model parameters, which is critical for practical applications. Additionally, we develop a new method to estimate tuning parameters for out-of-sample areas, enabling the derivation of both area-specific regression coefficients and sampling variances, even for out of sample areas.
Our simulation studies highlight the advantages of the proposed CLS method, which consistently outperforms the MR method in terms of both bias and RRMSPE, particularly when the assumptions of constant regression coefficients and sampling variances across areas are violated. These results underscore the effectiveness of our approach in handling heterogeneous and complex data structures. It is important to note that estimates for out-of-sample areas, whether derived from CLS or MR, remain synthetic due to the absence of direct observations. Future research could focus on reducing the synthetic nature of these estimates by generating pseudo-random effect estimates or incorporating other information, such as spatial correlations among areas.
The field of SAE is also evolving with the integration of non-traditional data sources, such as satellite imagery, mobile phone data, and geospatial information, alongside the use of machine learning methods to address model misspecification. For instance, Jean et al. (2016) and others illustrate how combining machine learning algorithms with SAE techniques can enhance poverty estimates, particularly in data-scarce regions. More recently, Merfeld et al. (2024) showed that integrating survey data with publicly available geospatial data using SAE can help overcome some limitations of survey data and provide more geographically disaggregated estimates of various SDG indicators at the subnational level. These advancements offer promising avenues for refining the estimation of poverty measures in contexts where traditional survey data alone are insufficient.

### v2

@@ 6.2 Model-based simulation II
Table 3: Performance of predictors of small area FGT measures for all areas when out-of-sample areas are involved.
Predictor Results (%) for the following scenarios
Headcount Ratio Poverty Gap
(0, 0) ($\beta$, 0) ($\beta$, $\sigma_{\epsilon}^{2}$) (0, 0) ($\beta$, 0) ($\beta$, $\sigma_{\epsilon}^{2}$)
Average Relative Bias
MR 8.300 26.301 26.304 13.398 25.533 26.092
MRE 8.036 25.838 25.835 13.040 24.994 25.531
CLS 7.143 6.459 6.315 10.944 8.732 8.623
CLSE 6.874 6.303 6.158 10.530 8.621 8.516
SELL 15.314 84.199 84.272 25.550 91.191 91.973
Average RRMSPE
MR 32.424 36.610 37.311 45.137 40.147 42.183
MRE 32.176 36.161 36.845 44.749 39.605 41.592
CLS 32.542 14.040 13.972 44.976 18.037 18.176
CLSE 32.279 13.883 13.815 44.529 17.919 18.026
SELL 54.209 91.022 91.338 77.59 102.120 103.529
Table 4: Performance of predictors of small area FGT measures for out-of-sample areas.
Predictor Results (%) for the following scenarios
Headcount Ratio Poverty Gap
(0, 0) ($\beta$, 0) ($\beta$, $\sigma_{\epsilon}^{2}$) (0, 0) ($\beta$, 0) ($\beta$, $\sigma_{\epsilon}^{2}$)
Average Relative Bias
MR 17.198 80.313 84.274 27.654 83.879 85.204
MRE 16.623 78.924 82.921 26.856 82.112 83.387
CLS 16.985 21.112 20.708 27.042 25.127 25.998
CLSE 16.332 20.730 20.326 26.090 24.699 25.568
SELL 18.496 215.009 218.421 29.732 237.951 240.449
Average RRMSPE
MR 57.993 97.731 101.802 82.577 104.957 109.189
MRE 57.396 96.426 100.200 81.673 103.165 107.287
CLS 57.842 31.468 31.643 82.088 40.510 43.643
CLSE 57.192 31.083 31.255 81.029 40.055 43.168
SELL 61.842 226.861 232.650 87.297 240.298 240.377
Overall, our findings should be interpreted within the scope of the data-generating mechanisms examined in the simulation study. Under heterogeneous data-generating scenarios, the results indicate that the proposed method outperforms the existing approaches. In contrast, when the traditional homogeneous NER assumptions are satisfied, the MR/MRE predictors remain competitive and may be preferable due to their more parsimonious structure.
@@ 7 Application: poverty mapping for Albania
Poverty mapping is a crucial tool for understanding the spatial distribution of poverty and guiding the implementation of poverty alleviation programs. In Albania, where the economy is predominantly rural and income is difficult to measure accurately, poverty is assessed using a consumption-based measure (Betti et al., 2003). In this section, we use 2002 Living Standards Measurement Survey (LSMS) data to estimate poverty indicators for 374 municipalities in Albania. To supplement this information, we incorporate auxiliary data from the 2001 Census, which covers 726,895 households across Albania. The descriptive statistics of these datasets have been presented in Section 3, providing an overview of their key characteristics and ensuring their suitability for small area estimation.
@@
Household consumption expenditure is used as the primary welfare variable, denoted by $E_{ij}$. We apply a logarithmic transformation to obtain $Y_{ij}$ as the response variable. The covariates include household size, dwelling facilities, and ownership variables, based on prior poverty assessment studies in Albania (Betti et al., 2003; Tzavidis et al., 2008). These covariates are described in detail in Section 3.
While these covariates were selected to ensure comparability with earlier studies, they are context-specific and may not be directly generalizable to other regions or applications. In general small area estimation (SAE) settings, model specification and covariate selection are critical components that can substantially influence predictive performance and statistical inference. Accordingly, conclusions regarding the relative performance of alternative predictors should be interpreted conditional on the adopted model specification. The poverty line is set at 4,891 Leks per month for all districts, following Betti et al. (2003).
In our implementation, the area-level covariates $\bar{\mathbf{Z}}_{i}$ entering the logit model for $\tau_{i}$ are computed from the 2001 Census, that is, they represent finite-population means for each municipality. When population-level aggregates are unavailable in other applications, $\bar{\mathbf{Z}}_{i}$ may be replaced by design-unbiased estimators derived from the survey data, with appropriate adjustments to account for the additional estimation variability. Alternatively, a two-stage modeling strategy or calibration to external aggregate sources may be employed. We discuss these practical alternatives and their implications for uncertainty assessment in the Appendix.
Although the 2002 Albania LSMS data were collected using a two-stage cluster sampling design, cluster identifiers are not available in the dataset used for this analysis. Consequently, we model municipalities as small areas and include area-level random effects only. With this specification, the covariance structure $\mathbf{V}_{i}$ captures dependence among units within areas through the area effect, but it does not explicitly model additional intra-cluster correlation arising from primary sampling units. If detailed cluster information were available, an extension of the NERHDP framework to incorporate cluster-within-area random effects would be a natural and practically relevant generalization, potentially improving variance estimation and predictive performance.
To examine the existence of heterogeneity across municipalities, we fitted separate regression models between $Y_{ij}$ and the covariates for each municipality. Figures 4 and 5 show boxplots of the estimated regression coefficients and standardized residuals, respectively. These results reveal potential variation across municipalities, providing empirical support for the use of a model that accommodates heterogeneity. Additional diagnostic analyses, including assessments of normality, identification of outliers and high-leverage points, and tests for heterogeneity in regression coefficients, further support the suitability of the NERHDP model for this application. Details of these diagnostics are provided in the Appendix. Consistent with these findings, the model-based simulation results in Section 6 further demonstrate that in the simulation setup, the CLS method under the NERHDP model performs well, showing (i) substantially lower bias and RRMSPE than the MR method when heterogeneity in regression coefficients and/or sampling variances exists, and (ii) improved computational efficiency relative to the algorithm proposed by Lahiri and Salvati (2023), making it more feasible for large-scale applications.
@@
Figure 4: Distributions of estimated regression coefficients fitted for each of the 213 municipalities of Albania.
Figure 5: Distributions of the standardized residuals by municipality.
@@
We apply the CLS method to obtain EBPs of HCRs and PGs by municipality, denoted as $\hat{F}_{0i}^{CLS}$ and $\hat{F}_{1i}^{CLS}$, respectively. The MSPEs of these FGT measures are estimated using the parametric bootstrap estimator $\text{mspe}_{*}(\hat{F}_{\alpha i}^{CLS})$ with $B=100$ replicates. We then estimate the associated coefficients of variation (CVs) using:
@@
Following Molina and Rao (2010), we also compute direct estimators of the FGT measures $\hat{F}_{\alpha i}^{w}$ and their variance estimates $\widehat{\text{Var}}_{D}(\hat{F}_{\alpha i}^{w})$ for sampled municipalities using weighted sample means (see Section 3).
To assess the quality of our CLS estimator, we use a set of diagnostic tools based on the requirement that model-based small area estimates should align with the corresponding survey-weighted direct estimates, albeit more precise (Marino et al., 2019). Correlation coefficients between CLS and direct estimates for the HCR and PG are 0.796 and 0.802, respectively, suggesting strong agreement between the two approaches. Based on heuristic arguments, we adopt the goodness-of-fit diagnostic proposed by Brown et al. (2001) to further evaluate coherence:
@@
With $W=119.59$ for HCR and $W=175.75$ for PG, both are below the 95th percentile of a $\chi^{2}$ distribution with 212 degrees of freedom ($\chi^{2}_{212,0.95}=177.39$), indicating that CLS estimates are coherent with direct estimates.
Table 5 presents selected results across municipalities with varying sample sizes. The CLS estimators consistently exhibit lower estimated CVs than those of the direct estimators, even in areas with small samples. In several municipalities where direct estimates are zero, estimated CVs are undefined, highlighting a limitation of the direct estimation approach.
Table 5: Population size, sample size, direct and CLS estimates of HCRs (%) and PGs (%), and the estimated CVs (%) of direct and CLS estimators for the Albania municipalities with sample size closest to minimum, first quartile, median, third quartile, and maximum.
HCR PG Municipality $N_{i}$ $n_{i}$ $\hat{F}_{0i}^{w}$ $\hat{F}_{0i}^{CLS}$ CV($\hat{F}_{0i}^{w}$) CV($\hat{F}_{0i}^{CLS}$) $\hat{F}_{1i}^{w}$ $\hat{F}_{1i}^{CLS}$ CV($\hat{F}_{1i}^{w}$) CV($\hat{F}_{1i}^{CLS}$) Hajmel 1111 6 33.33 23.82 56.96 27.01 1.52 5.92 70.56 31.31 Orenje 1419 16 12.50 20.69 66.19 15.06 2.00 4.97 66.03 17.44 B.Curri 1488 40 20.00 24.54 32.82 8.57 5.65 5.51 43.01 9.56 Gramsh 2503 64 23.55 17.97 24.48 0.69 4.75 3.64 38.93 21.90 Tirane 89764 600 12.43 11.72 10.91 4.46 2.49 2.31 14.71 4.72
Figure 6 compares the empirical cumulative distribution functions (ECDFs) of the estimated CVs for CLS and direct estimators. The ECDFs of the CLS estimators dominate those of the direct estimators for both HCR and PG, illustrating their improved precision. Notably, about 78% (HCR) and 89% (PG) of municipalities have direct estimate with estimated CVs exceeding the conventional 33% reliability threshold in small area estimation (Marino et al., 2019). In contrast, these percentages drop to 28% and 34% under the CLS approach.
@@
Figures 7 and 8 present municipality-level poverty maps for Albania using direct and CLS estimates, respectively. The white areas in Figure 7 represent regions where direct poverty estimates are unavailable due to zero sample size, underscoring a key limitation of the direct estimation approach in small-area analysis. In contrast, Figure 8 presents poverty maps generated using our proposed CLS method, which provides more comprehensive and reliable estimates across all municipalities. The CLS maps indicate that municipalities in the Bulqize district, such as Ostren, Gjorice, Shupenze, and Bulqize, exhibit the highest HCR and PG values, significantly exceeding those observed in other municipalities. Similarly, municipalities like Gjoaaj in the Peqin district, Katund I in the Durres district, and Arren in the Kukes district also show relatively high poverty levels, highlighting the need for targeted policy interventions to improve living conditions in these areas. In contrast, municipalities in the southern regions display substantially lower poverty levels, reflecting better economic conditions and higher living standards. Although no formal field validation study is available for these municipality-level SAE results, the spatial patterns obtained here are broadly consistent with earlier poverty mapping analyses for Albania (Betti et al., 2003; Tzavidis et al., 2008), providing indirect empirical support for their plausibility.
Figure 7: Municipality-level direct estimates of HCRs and PGs in Albania.
Figure 8: Municipality-level CLS estimates of HCRs and PGs in Albania.
Overall, the CLS method provides a more complete and reliable representation of poverty distribution in Albania. By producing estimates for all municipalities, it addresses the data limitations of direct estimation and enables a more comprehensive understanding of regional poverty disparities. The identification of high-poverty areas in the northern and central regions highlights the need for geographically targeted poverty alleviation programs, while the lower poverty levels in the south suggest regional inequalities that should inform national development strategies. The robustness of the CLS estimates, even in municipalities with small sample sizes, suggests that, when the data structure and underlying conditions are comparable to those considered in our simulation studies, our method may offer practical value for policy-oriented poverty mapping.
@@ 8 Discussion
The proposed model bridges the gap between existing approaches that either assume identical model parameters across areas, potentially leading to misspecification, or rely on random slope or sampling variance specifications that demand strong distributional assumptions. By leveraging area-specific estimating equations and appropriately constructed residuals, we offer a framework for deriving empirical best prediction of complex poverty measures that captures variability across areas without the instability associated with fixed effects in small samples. On the computational side, we introduce a novel, computationally efficient, and data-driven procedure for estimating model parameters, which is critical for practical applications. Additionally, we develop a new method to estimate tuning parameters for out-of-sample areas, enabling the derivation of both area-specific regression coefficients and sampling variances, even for out of sample areas.
Based on the Albania data, our simulation studies highlight the advantages of the proposed CLS method in prediction accuracy, which outperforms the MR method in terms of both bias and RRMSPE, when the assumptions of constant regression coefficients and/or sampling variances across areas are violated. These results demonstrate the effectiveness of our approach in handling heterogeneous and complex data structures. It is important to note, however, that no single SAE method is universally optimal. The relative performance of NERHDP, traditional NER-based EBP, or alternative approaches such as ELL depends on the underlying data structure, the degree of heterogeneity across areas, model specification, and the availability and quality of auxiliary information. The proposed method is particularly advantageous when regression coefficients and/or sampling variances vary across areas, but under homogeneous settings simpler models may perform equally well or better. We therefore encourage future research to further assess the methodology under a broader range of simulation scenarios.
It is important to note that estimates for out-of-sample areas, whether derived from CLS or MR, remain synthetic due to the absence of direct observations. Although the proposed method performs well in out-of-sample areas when strong auxiliary information is present, future research on reducing the synthetic nature of these estimates is warranted.
In this paper, we empirically evaluated the performance of the proposed model parameter estimators and small area predictors. Investigation of asymptotic properties will be an interesting future research topic, but for such research we would need a suitable theoretical framework that is likely to require large area specific sample sizes. We did not explore such asymptotic theory in this paper because in our applications the assumption of large area specific sample sizes is not realistic. Future work could also investigate formal inference procedures, more detailed structures for the tuning parameter $\tau_{i}$ and model selection strategies for the NERHDP model, since these components are essential for practical implementation.
While the proposed NERHDP model demonstrates promising results based on the Albania data in reducing bias and improving precision, its performance may vary under different data structures and sampling conditions. For instance, under complex survey design, such as multi-stage sampling conducted, with available cluster identifiers and rich contextual variables, alternative modeling strategies may better capture the data structure and yield improved performance.
In our simulation studies, EBP-type approaches tend to outperform the ELL-type predictor. However, in practice, the relative performance of EBP- and ELL-type methods depends critically on the underlying data structure, the availability of auxiliary information at different levels of aggregation, and the specification of the variance structure. As highlighted in recent review articles (e.g., Molina et al., 2022; Das and Chambers, 2024), neither class of methods is uniformly superior. ELL-type estimators may benefit from rich contextual information and flexible variance modeling, whereas EBP-type predictors may perform well when the nested error regression framework adequately represents the data-generating mechanism.
Open questions also remain regarding the estimation of variance components, particularly for random effects. Although no zero or negative variance estimates were observed in our study, the algorithm does not guarantee positivity of variance estimates of random effects. Non-positive variance estimates would compromise the bootstrap-based uncertainty measures proposed here. Further research is therefore needed to ensure valid variance estimation and enhance the robustness of the NERHDP framework.
Finally, similar to much of the existing literature on empirical best prediction in the poverty mapping literature, this paper assumes normality for both the random effects and the error terms. Nevertheless, we acknowledge that normality is not a prerequisite for the development of EBP methodology. Our empirical findings suggest that future research should explore alternative parametric distributions motivated by observed data characteristics, thereby improving model flexibility and robustness in practical applications.
Appendix: Diagnostic checks for Albanian application
While a comprehensive treatment of model diagnostics is beyond the scope of the current paper, we acknowledge the importance of assessing model adequacy in small area estimation contexts. We conducted several diagnostics and discussion to support the plausibility of the usage of NERHDP model in the Albanian application.
1. Normality of residuals:
To assess the plausibility of the normality assumption underlying the nested error regression framework, we fitted a two-level linear mixed model to the Albanian data and examined the empirical distributions of both area-level random effects and unit-level residuals. Normal probability plots are displayed in Figure 9. The Shapiro–Wilk tests yield p-values of $2.29\times 10^{-10}$ at level 1 and $2.78\times 10^{-6}$ at level 2, indicating departures from strict normality.
These results suggest that normality should be viewed as an approximation rather than a literal description of the data-generating process. Such deviations are not unexpected in welfare and expenditure data, even after logarithmic transformation. Within the superpopulation framework adopted here, normality is primarily a working modeling assumption that facilitates derivation of conditional distributions required for empirical best prediction.
Importantly, the proposed NERHDP approach incorporates robustness features through influence-function-based estimating equations (with Huber-type weights), which mitigate the impact of outliers and moderate departures from distributional assumptions. Moreover, empirical best predictors under nested error regression models are known to be relatively stable under mild deviations from normality, particularly when the primary goal is conditional prediction rather than exact likelihood-based inference.
Taken together, these diagnostics motivate the use of a flexible and robust modeling strategy rather than strict reliance on Gaussian assumptions. Future work may consider semiparametric or distributionally robust extensions of the conditional prediction framework to further relax Gaussian assumptions.
Figure 9: Normal probability plots of level 1 (left) and level 2 residuals (right) derived by fitting a two level linear mixed model to sample data.
2. Outliers and high leverage points:
Following Zewotir and Galpin (2007), we examined potential outliers and high leverage points in the sample data. Specifically, the diagonal elements $s_{ii}$ of the matrix $\mathbf{S}=\sigma^{2}_{\epsilon}\mathbf{P}$, with $\mathbf{P}=\mathbf{V}^{-1}-\mathbf{V}^{-1}\mathbf{X}(\mathbf{X}^{\prime}\mathbf{V}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\prime}\mathbf{V}^{-1}$, are used to identify high leverage points, where $\mathbf{X}$ is the design matrix and $\mathbf{V}$ is the covariance matrix of the linear mixed model. The diagnostic involves plotting $s_{ii}$ against $\hat{e}^{2}_{i}/\hat{\mathbf{e}}^{\prime}\hat{\mathbf{e}}$, where $\hat{e}_{i}$ denotes the EBLUP residual. Points are expected to cluster around the upper-left corner of the plot. Those in the lower-left corner (small $s_{ii}$) indicate high leverage, while those on the right (large relative squared residuals) suggest outliers. As shown in Figure 10, both outliers and high leverage points are potentially present.
Figure 10: Plot to detect both outliers and high leverage points: $s_{ii}$ versus $\hat{e}^{2}_{i}/\hat{\mathbf{e}}^{\prime}\hat{\mathbf{e}}$
These diagnostics highlight the need for a more robust and flexible approach. Accordingly, we fitted the NERHDP model to the Albanian data, setting the tuning constant in the Huber influence function to $c=1.345$.
3.
In addition, we conducted Hausman test (Hausman, 1978) to assess the heterogeneity of regression coefficients. The test statistic is defined as: $h_{i}=(\hat{\bm{\beta}}-\hat{\bm{\beta}}_{i})^{\prime}\bigg(\widehat{Var}(\hat{\bm{\beta}})-\widehat{Var}(\hat{\bm{\beta}}_{i})\bigg)^{-1}(\hat{\bm{\beta}}-\hat{\bm{\beta}}_{i})$ with $\widehat{Var}(\hat{\bm{\beta}}_{i})=(\sum_{l=1}^{m}\mathbf{X}_{l}\hat{\mathbf{V}}^{-1}_{l;i}\mathbf{X}_{l}^{\prime})^{-1}$. Under the null hypothesis that regression coefficients are homogeneous across small areas, this statistic approximately follows a $\chi^{2}_{(p-1)\times m}$ distribution. The observed value of 6991.664 yields a very small $p$-value, leading us to reject the null hypothesis. This provides strong evidence of heterogeneity in the regression coefficients and aligns with the boxplots in Figure 8, further supporting the use of the NERHDP model for the Albanian data.